An adult pass for a carnival costs $5 more than a children's pass. When 410 adult and 562 children's passes were sold, the total revenue was $6424. Find the cost of an adult pass.
Answer by macston(5194)

.
A=cost of adult pass: C=cost of child pass=A-$5
410A+562C=$6424 Substitute for C.
410A+562(A-$5)=$6424
410A+562A-$2810=$6424 Add $2810 to each side.
972A=$9234 Divide each side by 972.
A=$9.50 ANSWER: Adult passes cost $9.50.
CHECK:
C=A-$5=$9.50-$5=$4.50
410A+562C=$6424
410($9.50)+562($4.50)=$6424
$3895+$2529=$6424
$6424=$6424